A certain drug is used to treat asthma. In a clinical trial of the drug. 27 of 285 treated subjects experienced headaches (baplease show all work
Answer #1

Part a)

Left tailed test

Part b)

Test Statistic :-
Z = ( P - P0) / √(P0 * q0 / n)
Z = ( 0.1019 - 0.1 ) / √(( 0.1 * 0.9) /265)
Z = 0.10

Part c)

P value = P ( Z < 0.10 ) = 0.5408

Part d)

H0 :- P = 0.10

Reject null hypothesis if P value < α = 0.05
Since P value = 0.5408 > 0.05, hence we fail to reject the null hypothesis
Conclusion :- We Fail to Reject H0

Option B is correct

Part e)

Option B is correct